OpenStudy (anonymous):

PLS HELP?? Find the area of a sector with an arc length of 60 in. and a radius of 15 in.

OpenStudy (anonymous):

I tried drawing it out.

OpenStudy (nurali):

l = theta/360 * 2Pi * r 60 = theta/360 * 2Pi * 15 theta/360 = 2/Pi theta = 720/Pi Area = theta/360 * Pi r^2 Area = 720/Pi * 1/360 * Pi 15^2 Area = 2*225 = 450 ins^2

OpenStudy (anonymous):

thx!(:

OpenStudy (nurali):

My Pleasure.